OpenStudy (anonymous):

how to integrate {1/(kx^2-gm)} where g and m and k are constants

OpenStudy (just_one_last_goodbye):

This math is so above me knowledge I thought It was spam X'D

OpenStudy (freckles):

do you know if the constants are positive or not?

OpenStudy (anonymous):

they are positive

OpenStudy (freckles):

awesome use partial fractions

OpenStudy (freckles):

\[\frac{A}{\sqrt{k}x-\sqrt{gm}}+\frac{B}{\sqrt{k}x+\sqrt{gm}}=\frac{1}{kx^2-gm}\]

OpenStudy (anonymous):

i think i can take it from here, thank you for the great hint !

OpenStudy (freckles):

tell me what you find for A and B.

OpenStudy (freckles):

you will get something in terms of g and m

OpenStudy (kainui):

Just curious, where did this problem come from? It looks like it might have to do with solving a differential equation involving a spring and gravity or something interesting.

OpenStudy (anonymous):

True, i am modelling throwing a basketball, and trying to solve for the speed, this came up when i am analyzing the vertical motion of the ball